OpenStudy (anonymous):

x-z=5 y+3z=12 2x+y=7 with steps please its kinda confusing me

OpenStudy (mathstudent55):

Take the third equation and solve for y. Can you do that?

OpenStudy (anonymous):

show me how to doit.. please

OpenStudy (mathstudent55):

The third equation is: 2x + y = 7 To solve for y, we need to get y by itself. That means we need to subtract 2x from both sides.

OpenStudy (anonymous):

just show me how to solve for y after that i can do it

OpenStudy (mathstudent55):

2x + y = 7 -2x -2x ------------------ y = 7 - 2x This is the third equation solved for y. Now substitute 7 - 2x for y in the second equation.

OpenStudy (anonymous):

oh baby i love you i got it like a boss... THX i will be your fan for you now!!

OpenStudy (mathstudent55):

Once you do the substitution above, then you can solve that new equation together with the first one as a system of two equations in 2 unknowns, x and z.

OpenStudy (mathstudent55):

You're welcome.